You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100414 (14 Apr 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 April13-14

Observer: K.Boutsia, R. Speziali
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O. Kuhn
Instrument: LUCIFER1

Summary

Still a windy night. The dome opened ~1.5hr after twilight.
During the whole night the GCS and the DIMM were giving discrepant results. The seeing measured by imexam on some acquisition images was <1.0.
(08:00 UT) Time lost for an error in OFFSET TELESCOPE command.
(08:50 UT) Shortly after the dome was closed down due to increasing wind speed.

Data Summary

Project Objects Filter ExpTime Notes
Mannucci HDF-N 210zHJK + H 2hr arc+flat+telluric done

Details

3:20 - Dome Open - Pointing correction procedure using ACT0305 (high-el) ACT0085 (low-el)

3:37 - End of pointing correction procedure

3:38 - Pointing refinement using Perssons BS9145 - ptcorr

3:44 - collimating on BS9145

3:49 - collimation ok - seeing 1"

3:50 - Preset for field acquisition - various images for field recognition

4:00 - MOS ACQ

4:03 - slit image

4:05 - End of field acquisition procedure *

4:06 - Start HDF script - 1h in H band - Seeing (0.8GCS - 1.2 DIMM ?!)

4:52 - GUIDE STAR LOST! 1.5" dimm

5:14 - End of the script

5:15 - Through slit image to measure the seeing - DIMM reports 1.6"!

5:20 - Image Acquisition to measure the seeing (0.8 arcsec) - GCS and DIMM not reliable.

5:30 - Start the second HDF script - 1h in H band

6:38 - End of the script

6:39 - Start Calibration - calibration unit in for arc e flat

6:39 - Arc script

6:45 - Flat script

6:50 - Slew to tellurinc star. No star!

6:55 - Execute ptcorr on the telluric

7:00 - Telluric integration script

7:06 - End of Calibration

7:10 - Slew to HDF science field

7:15 - Field acquisition #1

7:30 - guide star lost after applying an offset!! Telescope error

7:35 - Telescope jumps on the following image

7:36 - Pointing correction WT10_385

7:42 - Field acquisition #2!!

7:49 - End of acquisition procedure

7:50 - start HDF script 1h in H (seeing measured on image= 0.85)

8:05 - ERROR on the 3rt OFFSET - Telescope Error

8:26 - Start acquisition #3!!!

8:30 - Abort acquisition

8:31 - Changing field since the guide star is too faint. Slewing to Q1422

8:40 - Not converging since we are facing the wind!

8:48 - Closing Dome due to strong wind
Topic revision: r7 - 14 Apr 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback